OpenStudy (anonymous):

Let's solve something simple and fun! Solve the amount of pennies and the amount of nickels. Hints: $0.17 total amount, 9 coins p+n=9 p+5n=17

OpenStudy (anonymous):

p + n = 9 --> p = 9 - n p + 5n = 17 9 - n + 5n = 17 9 + 4n = 17 4n = 17 - 9 4n = 8 n = 2 p + n = 9 p + 2 = 9 p = 9 - 2 p = 7 check... p + 5n = 17 7 + 5(2) = 17 7 + 10 = 17 17 = 17 (correct) 7 pennies and 2 nickels